Calculate the molarity of a solution that contains 29.6 g AI2 (SO4)3 in 238.7 mL water
1
Expert's answer
2020-04-23T13:06:24-0400

The molarity of the solution can be calculated using the following formula:

M = n / V = m / (Mr × V)

where M - molarity, n - number of moles, m - mass, Mr - molecular weight, V - volume.

As Mr(Al2(SO4)3) = 342 g/mol and V = 238.7 mL = 0.2387 L

M = 29.6 g / (342 g/mol × 0.2387 L) = 0.363 mol/L = 0.363 M


Answer: 0.363 M